OpenStudy (anonymous):

Billy started with 2 poems in his journal. Then he started writing 4 poems each day. Which of the following graph represents Billy's poem writing?

OpenStudy (anonymous):

@mathstudent55 @freckles

OpenStudy (mathstudent55):

Before we look at graphs, can you come up with an equation for this function? Let's say the number of days is x, and the number of poems is y.

OpenStudy (mathstudent55):

the number of poems = a function of the number of days y = a function of x

OpenStudy (anonymous):

so the slope could be 2?

OpenStudy (mathstudent55):

You can use the slope-intercept form: \(y = mx + b\) b is the y-intercept. That is the number you have at zero x-value. The slope is what accounts for what changes, in this case, daily.

OpenStudy (anonymous):

so y=2x+1

OpenStudy (mathstudent55):

the number of poems = a function of the number of days y = a function of x y = fixed beginning amount + daily additional amount

OpenStudy (mathstudent55):

No. Look at my last line above. y, the number of poems, is the sum of the beginning amount and the daily additional amount. What is the beginning amount? Hint: "Billy started with 2 poems in his journal."

OpenStudy (anonymous):

2x

OpenStudy (anonymous):

would the points be (4,2) and (0,2)

OpenStudy (mathstudent55):

No. No x. The fixed amount never changes. It is not dependent on x. The fixed amount is simply 2 since Billy started with 2 poems in his journal.

OpenStudy (anonymous):

Ok, so the points would be (0,2) (1,6) and (2,10)

OpenStudy (anonymous):

am i right?

OpenStudy (mathstudent55):

The daily additional amount depends on the day. He adds 4 poems per day. Day is x, so after day 1 he has 4 * 1 extra poems. after day 2 he has 4 * 2 extra poems after day 3 he has 4 * 3 extra poems after day 4 he has 4 * 4 extra poems this pattern goes on and on... after day x he has 4 * x extra poems

OpenStudy (mathstudent55):

Since y, the number of poems in his journal is given by the fixed number of poems he started with plus the daily additional poems, we can write: y = 2 + 4x which can be rewritten as y = 4x + 2

OpenStudy (anonymous):

Ok! Thanks :)

OpenStudy (anonymous):

what would be some points if I graphed that?

OpenStudy (mathstudent55):

Yes, your points are correct because they are all points in the equation y = 4x + 2

OpenStudy (anonymous):

(0,2) (1,6) ??

OpenStudy (anonymous):

If I open another can u help again?

OpenStudy (mathstudent55):

sure